OpenStudy (anonymous):

use a half-angle formula to find the exact value of each expression sin 105

OpenStudy (anonymous):

@aum

OpenStudy (aum):

105 = 90 + 15 sin(105) = sin(90+15) = sin(90)cos(15) + cos(90)sin(15) = 1*cos(15) + 0*sin(15) = cos(15).

OpenStudy (anonymous):

wait wouldnt it be 105*2

OpenStudy (aum):

\[ \cos^2(\frac a2) = \frac 12\{1+\cos(a) \} \]

OpenStudy (aum):

I am splitting 105 degrees into two parts: 90 degrees + 15 degrees. 90 degrees is a standard angle for which the sine and cosine are known values. So I use the sin(A+B) formula first: sin(A+B) = sin(A)cos(B) + cos(A)sin(B) with A = 90 and B = 15 degrees.

OpenStudy (aum):

In my first reply I showed that sin(105) = cos(15). Now we use the half angle formula (in my second reply) and find cos(15). Set a = 30 in the half angle formula and find cos(15) and that will be the same as sin(105).

OpenStudy (anonymous):

okay go on i get it

OpenStudy (anonymous):

what should i do next

OpenStudy (aum):

same thing we did for sin(15) in the previous problem. set a = 30 in the half angle formula above and find cos(15).

OpenStudy (anonymous):

okay thank you

OpenStudy (aum):

you are welcome.

OpenStudy (anonymous):

one last one?